OpenStudy (anonymous):

The variables x and y vary directly. Write an equation that relates x and y, when x = 4 and y = 2.

OpenStudy (nali):

since x and y vary directly that means that the equaiton will be similar to this form y=kx where k is the constant or y=x/k Not y=k/x >>> which means that y varies INDIRECTLY with x

OpenStudy (nali):

so what is the relationship betwee the y=2 and x =4 y=x/k 2=4/k so k=2 meaning y=x/2 >>> the equaiton